40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log, Page 5 of 344  Not logged in ELOG logo
ID Date Author Type Category Subject
  17119   Wed Aug 31 01:30:53 2022 KojiUpdateGeneralAlong the X arm part 4

Behind the X arm tube

Attachment 1: PXL_20220831_020757504.jpg
PXL_20220831_020757504.jpg
Attachment 2: PXL_20220831_020825338.jpg
PXL_20220831_020825338.jpg
Attachment 3: PXL_20220831_020856676.jpg
PXL_20220831_020856676.jpg
Attachment 4: PXL_20220831_020934968.jpg
PXL_20220831_020934968.jpg
Attachment 5: PXL_20220831_021030215.jpg
PXL_20220831_021030215.jpg
  17118   Wed Aug 31 01:25:37 2022 KojiUpdateGeneralAlong the X arm part 3

 

 

Attachment 1: PXL_20220831_020455209.jpg
PXL_20220831_020455209.jpg
Attachment 2: PXL_20220831_020534639.jpg
PXL_20220831_020534639.jpg
Attachment 3: PXL_20220831_020556512.jpg
PXL_20220831_020556512.jpg
Attachment 4: PXL_20220831_020606964.jpg
PXL_20220831_020606964.jpg
Attachment 5: PXL_20220831_020615854.jpg
PXL_20220831_020615854.jpg
Attachment 6: PXL_20220831_020623018.jpg
PXL_20220831_020623018.jpg
Attachment 7: PXL_20220831_020640973.jpg
PXL_20220831_020640973.jpg
Attachment 8: PXL_20220831_020654579.jpg
PXL_20220831_020654579.jpg
Attachment 9: PXL_20220831_020712893.jpg
PXL_20220831_020712893.jpg
  17117   Wed Aug 31 01:24:48 2022 KojiUpdateGeneralAlong the X arm part 2

 

 

Attachment 1: PXL_20220831_020307235.jpg
PXL_20220831_020307235.jpg
Attachment 2: PXL_20220831_020333966.jpg
PXL_20220831_020333966.jpg
Attachment 3: PXL_20220831_020349163.jpg
PXL_20220831_020349163.jpg
Attachment 4: PXL_20220831_020355496.jpg
PXL_20220831_020355496.jpg
Attachment 5: PXL_20220831_020402798.jpg
PXL_20220831_020402798.jpg
Attachment 6: PXL_20220831_020411566.jpg
PXL_20220831_020411566.jpg
Attachment 7: PXL_20220831_020419923.jpg
PXL_20220831_020419923.jpg
Attachment 8: PXL_20220831_020439160.jpg
PXL_20220831_020439160.jpg
Attachment 9: PXL_20220831_020447841.jpg
PXL_20220831_020447841.jpg
  17116   Wed Aug 31 01:22:01 2022 KojiUpdateGeneralAlong the X arm part 1

 

Attachment 5: RF delay line was accommodated in 1X3B. (KA)

Attachment 1: PXL_20220831_015945610.jpg
PXL_20220831_015945610.jpg
Attachment 2: PXL_20220831_020024783.jpg
PXL_20220831_020024783.jpg
Attachment 3: PXL_20220831_020039366.jpg
PXL_20220831_020039366.jpg
Attachment 4: PXL_20220831_020058066.jpg
PXL_20220831_020058066.jpg
Attachment 5: PXL_20220831_020108313.jpg
PXL_20220831_020108313.jpg
Attachment 6: PXL_20220831_020131546.jpg
PXL_20220831_020131546.jpg
Attachment 7: PXL_20220831_020145029.jpg
PXL_20220831_020145029.jpg
Attachment 8: PXL_20220831_020203254.jpg
PXL_20220831_020203254.jpg
Attachment 9: PXL_20220831_020217229.jpg
PXL_20220831_020217229.jpg
  17115   Wed Aug 31 00:46:56 2022 KojiUpdateGeneralVertex Lab area to be cleaned

As marked up in the photos.

 

Attachment 5: The electronics units removed. Cleaning half way down. (KA)

Attachment 6: Moved most of the units to 1X3B rack ELOG 17125 (KA)

Attachment 1: PXL_20220831_015758208.jpg
PXL_20220831_015758208.jpg
Attachment 2: PXL_20220831_015805113.jpg
PXL_20220831_015805113.jpg
Attachment 3: PXL_20220831_015816628.jpg
PXL_20220831_015816628.jpg
Attachment 4: PXL_20220831_015826533.jpg
PXL_20220831_015826533.jpg
Attachment 5: PXL_20220831_015844401.jpg
PXL_20220831_015844401.jpg
Attachment 6: PXL_20220831_015854055.jpg
PXL_20220831_015854055.jpg
Attachment 7: PXL_20220831_015903704.jpg
PXL_20220831_015903704.jpg
Attachment 8: PXL_20220831_015919376.jpg
PXL_20220831_015919376.jpg
Attachment 9: PXL_20220831_021708873.jpg
PXL_20220831_021708873.jpg
  17114   Wed Aug 31 00:32:00 2022 KojiUpdateGeneralSOS and other stuff in the clean room

Salvage these (and any other things). Wrap and double-pack nicely. Put the labels. Store them and record the location. Tell JC the location.

Attachment 1: PXL_20220831_015137480.jpg
PXL_20220831_015137480.jpg
Attachment 2: PXL_20220831_015203941.jpg
PXL_20220831_015203941.jpg
Attachment 3: PXL_20220831_015230288.jpg
PXL_20220831_015230288.jpg
Attachment 4: PXL_20220831_015249451.jpg
PXL_20220831_015249451.jpg
Attachment 5: PXL_20220831_015424383.jpg
PXL_20220831_015424383.jpg
Attachment 6: PXL_20220831_015433113.jpg
PXL_20220831_015433113.jpg
  17113   Tue Aug 30 15:21:27 2022 TegaUpdateComputers3 FEs from LHO got delivered today

[Tega, JC]

We received the remaining 3 front-ends from LHO today. They each have a timing card and an OSS host adapter card installed. We also receive 3 dolphin DX cards. As with the previous packages from LLO, each box contains a rack mounting kit for the supermicro machine.

Attachment 1: IMG_20220830_144925325.jpg
IMG_20220830_144925325.jpg
Attachment 2: IMG_20220830_142307495.jpg
IMG_20220830_142307495.jpg
Attachment 3: IMG_20220830_143059443.jpg
IMG_20220830_143059443.jpg
  17112   Mon Aug 29 18:25:12 2022 CiciUpdateGeneralTaking finer measurements of the actuator transfer function

Took finer measurements of the x-arm aux laser actuator tranfer function (10 kHz - 1 MHz, 1024 pts/decade) using the Moku.

--------------------------------------

I took finer measurements using the moku by splitting the measurement into 4 sections (10 - 32 (~10^4.5) kHz, 32 - 100 kHz, 100 - 320 kHz, 320 - 1000 kHz) and then grouping them together. I took 25 measurements of each ( + a bonus in case my counting was off), plotted them in the attached notebook, and calculated/plotted the standard deviation of the magnitude (normalized for DC offset). Could not upload to the ELOG as .pdf, but the pdf's are in the .zip file.

--------------------------------------

Next steps are to do the same stdev calculation for phase, which shouldn't take long, and to use the vectfit of this better data to create a PZT inversion filter.

Attachment 1: PZT_TF_fine.png
PZT_TF_fine.png
Attachment 2: PZT_TF_fine_mag_stdev.png
PZT_TF_fine_mag_stdev.png
Attachment 3: ATF_fine.zip
  17111   Mon Aug 29 15:15:46 2022 TegaUpdateComputers3 FEs from LLO got delivered today

[JC, Tega]

We got the 3 front-ends from LLO today. The contents of each box are:

  1. FE machine
  2. OSS adapter card for connecting to I/O chassis
  3. PCI riser cards (x2)
  4. Timing Card and cable
  5. Power cables, mounting brackets and accompanying screws
Attachment 1: IMG_20220829_145533452.jpg
IMG_20220829_145533452.jpg
Attachment 2: IMG_20220829_144801365.jpg
IMG_20220829_144801365.jpg
  17110   Mon Aug 29 13:33:09 2022 JCUpdateGeneralLab Cleanup

The machine shop looked a mess this morning, so I cleaned it up. All power tools are now placed in the drawers in the machine shop. Let me know if there are any questions of where anything here is placed. 

Attachment 1: EDE63209-D556-41F1-9BF2-89CD78E3D7B7.jpeg
EDE63209-D556-41F1-9BF2-89CD78E3D7B7.jpeg
  17109   Sun Aug 28 23:14:22 2022 JamieUpdateComputersrack reshuffle proposal for CDS upgrade

@tega This looks great, thank you for putting this together.  The rack drawing in particular is great.  Two notes:

  1. In "1X6 - proposed" I would move the "PEM AA + ADC Adapter" down lower in the rack, maybe where "Old FB + JetStor" are, after removing those units since they're no longer needed.  That would keep all the timing stuff together at the top without any other random stuff in between them.  If we can't yet remove Old FB and the JetStor then I would move the VME GPS/Timing chassis up a couple units to make room for the PEM module between the VME chassis and FB1.
  2. We'll eventually want to move FB1 and Megatron into 1X7, since it seems like there will be room there.  That will put all the computers into one rack, which will be very nice.  FB1 should also be on the KVM switch as well.

I think most of this work can be done with very little downtime.

  17108   Fri Aug 26 14:05:09 2022 TegaUpdateComputersrack reshuffle proposal for CDS upgrade

[Tega, Jamie]

Here is a proposal for what we would like to do in terms of reshuffling a few rack-mounted equipments for the CDS upgrade. 

  • Frequency Distribution Amp - Move the unit from 1X7 to 1X6 without disconnecting the attached cables. Then disconnect power and signal cables one at a time to enable optimum rerouting for strain relief and declutter.  

 

  • GPS Receiver Tempus LX - Move the unit from 1X7 to 1X6 without disconnecting the attached cables. Then disconnect power and signal cables one at a time to enable optimum rerouting for strain relief and declutter.

 

  • PEM & ADC Adapter - Move the unit from 1X7 to 1X6 without disconnecting the attached cables. Disconnect the single signal cable from the rear of the ADC adapter to allow for optimum rerouting for strain relief.

 

  • Martian Network Switch - Make a note of all connections, disconnect them, move the switch to 1X7 and reconnect ethernet cables. 

 

  • MARTIAN NETWORK SWITCH CONNECTIONS
    # LABEL # LABEL
    1 Tempus LX (yellow,unlabeled) 13 FB1
    2 1Y6 HUB 14 FB
    3 C0DCU1 15 NODUS
    4 C1PEM1 16  
    5 RFM-BYPASS 17 CHIARA
    6 MEGATRON/PROCYON 18  
    7 MEGATRON 19 CISC/C1SOSVME
    8 BR40M 20 C1TESTSTAND [blue/unlabelled]
    9 C1DSCL1EPICS0 21 JetStar [blue/unlabelled]
    10 OP340M 22 C1SUS [purple]
    11 C1DCUEPICS 23 unknown [88/purple/goes to top-back rail]
    12 C1ASS 24 unknown [stonewall/yellow/goes to top-front rail]

     

I believe all of this can be done in one go followed by CDS validation. Please comment so we can improve the plan. Should we move FB1 to 1X7 and remove old FB & JetStor during this work?

Attachment 1: Reshuffling proposal

Attachment 2: Front of 1X7 Rack

Attachment 3: Rear of 1X7 Rack

Attachment 4: Front of 1X6 Rack

Attachment 5: Rear of 1X6 Rack

Attachment 6: Martian switch connections

Attachment 1: rack_change_proposal.pdf
rack_change_proposal.pdf
Attachment 2: IMG_20220826_131331042.jpg
IMG_20220826_131331042.jpg
Attachment 3: IMG_20220826_131153172.jpg
IMG_20220826_131153172.jpg
Attachment 4: IMG_20220826_131428125.jpg
IMG_20220826_131428125.jpg
Attachment 5: IMG_20220826_131543818.jpg
IMG_20220826_131543818.jpg
Attachment 6: IMG_20220826_142620810.jpg
IMG_20220826_142620810.jpg
  17107   Fri Aug 26 12:46:07 2022 CiciUpdateGeneralProgress on fitting PZT resonances

Here is an update of how fitting the resonances is going - I've been modifying parameters by hand and seeing the effect on the fit. Still a work in progress. Magnitude is fitting pretty well, phase is very confusing. Attempted vectfit again but I can't constrain the number of poles and zeros with the code I have and I still get a nonsensical output with 20 poles and 20 zeros. Here is a plot with my fit so far, and a zip file with my moku data of the resonances and the code I'm using to plot.

Attachment 1: PZT_fit.zip
Attachment 2: AUX_PZT_Actuator_narrow_fit_1.pdf
AUX_PZT_Actuator_narrow_fit_1.pdf
  17106   Thu Aug 25 16:39:31 2022 CiciUpdateGeneralI have learned the absolute basics of github

I have now added code/data to my github repository. (it's the little victories)

  17105   Thu Aug 25 16:05:51 2022 YehonathanUpdateSUSTrying to fix some SUS

I tried to lock the Y/X arms to take some noise budget. However, we noticed that TRX/Y were oscillating coherently together (by tens of percent), meaning some input optics, essentially PR2/3 are swinging. There was no way I could do noise budgeting in this situation.

I set out to debug these optics. First, I notice side motion of PR2 is very weakly damped .

The gain of the side damping loop (C1:SUS-PR2_SUSSIDE_GAIN) was increased from 10 to 150 which seem to have fixed the issue. Attachment 1 shows the current step response of  the PR2 DOFs. The residual Qs look good but there is still some cross-couplings, especially when kicking POS. Need to do some balancing there.

PR3 fixing was less successful in the beginning. I increased the following gains:

C1:SUS-PR3_SUSPOS_GAIN: 0.5 -> 30

C1:SUS-PR3_SUSPIT_GAIN: 3 -> 30

C1:SUS-PR3_SUSYAW_GAIN: 1 -> 30

C1:SUS-PR3_SUSSIDE_GAIN: 10 -> 50

But the residual Q was still > 10. Then I checked the input matrix and noticed that UL->PIT is -0.18 while UR->PIT is 0.39. I changed UL->PIT (C1:SUS-PR3_INMATRIX_2_1) to +0.18. Now the Q became 7. I continue optimizing the gains.

Was able to increase C1:SUS-PR3_SUSSIDE_GAIN: 50 -> 100.

Attachment 2 shows the step response of PR3. The change of the entry of the input matrix was very ad-hoc, it would probably be good to run a systematic tuning. I have to leave now, but the IFO is in a very misaligned state. PR3/2 should be moved to bring it back.

Attachment 1: PR2_Step_Response_Test_2022-08-25_16-23.pdf
PR2_Step_Response_Test_2022-08-25_16-23.pdf PR2_Step_Response_Test_2022-08-25_16-23.pdf PR2_Step_Response_Test_2022-08-25_16-23.pdf PR2_Step_Response_Test_2022-08-25_16-23.pdf
Attachment 2: PR3_Step_Response_Test_2022-08-25_18-37.pdf
PR3_Step_Response_Test_2022-08-25_18-37.pdf PR3_Step_Response_Test_2022-08-25_18-37.pdf PR3_Step_Response_Test_2022-08-25_18-37.pdf PR3_Step_Response_Test_2022-08-25_18-37.pdf
  17104   Thu Aug 25 15:24:06 2022 PacoHowToElectronicsRFSoC 2x2 board -- fandango

[Paco, Chris Stoughton, Leo -- remote]

This morning Chris came over to the 40m lab to help us get the RFSoC board going. After checking out our setup, we decided to do a very basic series of checks to see if we can at least get the ADCs to run coherently (independent of the DACs). For this I borrowed the Marconi 2023B from inside the lab and set its output to 1.137 GHz, 0 dBm. Then, I plugged it into the ADC1 and just ran the usual spectrum analyzer notebook on the rfsoc jupyter lab server. Attachment #1 - 2 shows the screen captured PSDs for ADCs 0 and 1 respectively with the 1137 MHz peaks alright.

The fast ADCs are indeed reading our input signals.


Before this simple test, we actually reached out to Leo over at Fermilab for some remote assistance on building up our minimally working firmware. For this, Chris started a new vivado project on his laptop, and realized the rfsoc 2x2 board files are not included in it by default. In order to add them, we had to go into Tools, Settings and add the 2020.1 Vivado Xilinx shop board repository path to the rfsoc2x2 v1.1 files. After a little bit of struggling, uninstalling, reinstalling them, and restarting Vivado, we managed to get into the actual overlay design. In there, with Leo's assistance, we dropped the Zynq MPSoC core (this includes the main interface drivers for the rfsoc 2x2 board). We then dropped an rf converter IP block, which we customized to use the right PLL settings. The settings, from the System Clocking tab were changed to have a 409.6 MHz Reference Clock (default was 122.88 MHz). This was not straightforward, as the default sampling rate of 2.00 GSPS was not integer-related so we had to also update that to 4.096 GSPS. Then, we saw that the max available Clock Out option was 256 MHz (we need to be >= 409.6 MHz), so Leo suggested we dropped a Clocking Wizard block to provide a 512 MHz clock input for the rfdc. The final settings are captured in Attachment # 3. The Clocking Wizard was added, and configured on its Output Clocks tab to provide a Requested Output Freq of 512 MHz. The finall settings of the Clocking wizard are captured in Attachment #4. Finally, we connected the blocks as shown in Attachment #5.

We will continue with this design tomorrow.

Attachment 1: adc0_1137MHz.png
adc0_1137MHz.png
Attachment 2: adc1_1137MHz.png
adc1_1137MHz.png
Attachment 3: rfdc_PLLsettings.png
rfdc_PLLsettings.png
Attachment 4: clockingwiz_settings.png
clockingwiz_settings.png
Attachment 5: blockIPdiag.png
blockIPdiag.png
  17103   Wed Aug 24 16:37:52 2022 CiciUpdateGeneralMore DFD/AUX PZT resonance measurements

Some more measurements of the PZT resonances (now zoomed in!) I'm adjusting parameters on our model to try and fit to it by hand a bit, definitely still needs improvements but not bad for a 2-pole 2-zero fit for now. I don't have a way to get coherence data from the moku yet but I've got a variety of measurements and will hopefully use the standard deviation to try and find a good error prediction...

 

Attachment 1: AUX_PZT_Actuator_narrow_fit.pdf
AUX_PZT_Actuator_narrow_fit.pdf
  17102   Wed Aug 24 12:02:24 2022 PacoUpdateSUSITMX SUS is sus UL glitches?

[Yehonathan, Paco]

This morning, while attempting to align the IFO to continue with noise-budgeting, we noted the XARM lock was not stable and showed glitches in the C1:LSC-TRX_OUT (arm cavity transmission). Inspecting the SUS screens, we found the ULSEN rms ~ 6 times higher than the other coils so we opened an ndscope with the four face OSEM signals and overlay the XARM transmission. We immediately noticed the ULSEN input is noisy, jumping around randomly and where bigger glitches correlated with the arm cavity transmission glitches. This is appreciated in Attachment #1.


Signal chain investigation

We'll do a full signal investigation on ITMX SUS electronics to try and narrow down the issue, but it seems the glitches come and go... Is this from the gold satamp box? ...

Attachment 1: ITMX_UL_badness_08242022.png
ITMX_UL_badness_08242022.png
  17101   Wed Aug 24 10:49:43 2022 CiciUpdateGeneralMeasuring DFD output/X-arm laser PZT TF with Moku

We measured the TF of the X-arm laser PZT using the Moku so we can begin fitting to that data and hopefully creating a digital filter to cancel out PZT resonances. 

-------------------------------------------------------------

We calculated the DFD calibration (V/Hz) using:

Vrf = 0.158 mV (-6 dBm), Km = 1 (K_phi = Km*Vrf), cable length = 45m,  Tau = cable length/(0.67*3*10^8 m/s) ~ 220 ns. 

We've taken some preliminary data and can see the resonances around 200-300 kHz.

---------------------------------------------------------

Next steps are taking more data around the resonances specifically, calibrating the data using the DFD calibration we calculated, and adjusting parameters in our model so we can model the TF.

 

Attachment 1: AUX_PZT_Actuator_nofit.pdf
AUX_PZT_Actuator_nofit.pdf
  17100   Tue Aug 23 22:30:24 2022 TegaUpdateComputersc1teststand OS upgrade - I

[JC, Tega, Chris]

After moving the test stand front-ends, chiara (name server) and fb1 (boot server) to the new rack behind 1X7, we powered everything up and checked that we can reach c1teststand via pianosa and that the front-ends are still able to boot from fb1. After confirming these tests, we decided to start the software upgrade to debian 10. We installed buster on fb1 and are now in the process of setting up diskless boot. I have been looking around for cds instructions on how to do this and I found the CdsFrontEndDebian10page which contains most of the info we require. The page suggests that it may be cleaner to start the debian10 installation on a front-end that is connected to an I/O chassis with at least 1 ADC and 1 DAC card, then move the installation disk to the boot server and continue from there, so I moved the disk from fb1 to one of the front-ends but I had trouble getting it to boot. I decided to do a clean install on another disk on the c1lsc front-end which has a host adapter card that can be connected to the c1bhd I/O chassis. We can then mount this disk on fb1 and use it to setup the diskless boot OS.

  17099   Tue Aug 23 14:59:15 2022 JCUpdateToolsNew Toolbox at Y-End

A new tool box has been placed at the Y-end! Each drawer has its label so PLEASE put the tools back in their correct location. In addition to this, Each tool has its assigned tool box, so PLEASE RETURN all tools to their designated tool box. The tools can be distinguished by a writing or heat shrink which corresponds to the color of the tool chest or location. Photo #2 is an example of how the tools have been marked.

Each toolbox from now on will contain a drawer for the folllowing: Measurements, Allen Keys, Pliers and Cutters, Screwdrivers, Zipties and Tapes, Allen Ball Drivers, Crescent Wrenches, Clamps, and Torque Wrenches/ Ratchets.

Attachment 1: 9AFD3E49-0C5B-4626-889A-0A5C62590AD7.jpeg
9AFD3E49-0C5B-4626-889A-0A5C62590AD7.jpeg
Attachment 2: 99EC2EB1-EEA0-4AD8-B6D7-A494431E91E5.jpeg
99EC2EB1-EEA0-4AD8-B6D7-A494431E91E5.jpeg
  17098   Mon Aug 22 19:02:15 2022 TegaUpdateComputersc1teststand rack mounting for CDS upgrade II

[Tega, JC]

Moved the rack to the location of the test stand just behind 1X7 and plan to remove the other two small test stand racks to create some space there.  We then mounted the c1bhd I/O chassis and 4 front-end machines on the test stand (see attachment 1).

Installed the dolphin IX cards on all 4 front-end machines: c1bhd, c1ioo, c1sus, c1lsc. I also removed the dolphin DX card that was previously installed on c1bhd.

Found a single OneStop host card with a mini PCI slot mounting plate in a storage box (see attachment 2). Since this only fits into the dual PCI riser card slot on c1bhd, I swapped out the full-length PCI slot OneStop host card on c1bhd and installed it on c1lsc, (see attachments 3 & 4).

 

Attachment 1: IMG_20220822_185437763.jpg
IMG_20220822_185437763.jpg
Attachment 2: IMG_20220822_131340214.jpg
IMG_20220822_131340214.jpg
Attachment 3: c1bhd.jpeg
c1bhd.jpeg
Attachment 4: c1lsc.jpeg
c1lsc.jpeg
  17097   Mon Aug 22 14:36:49 2022 ranaUpdateSUSOpen loop transfer function measurements for local damping loops of Core optics

for damping and OL loops, we typically don't measure the TF like this because it takes forever and we don't need that detailed info for anything. Just use the step responses in the way we discussed at the meeting 2 weeks ago. There's multiple elog entries from me and others illustrating this. The measurement time is then only ~30 sec per optic, and you also get the cross-coupling for free. No need for test-point channels and overloading, just use the existing DQ channels and read back the response from the frames after the excitations are completed.

  17096   Sat Aug 20 20:26:10 2022 AnchalUpdateSUSOpen loop transfer function measurements for local damping loops of Core optics

I made measurements of old optics OLTF today. I have reduced the file sizes of the plots and data now. It is interesting that it is allowed to read 9 channels simultaneously from c1mcs or c1sus models, even together. The situation with c1su2 is a bit unclear. I was earlier able to take measurements of 6 channels at once from c1su2 but not I can't read more than 1 channel simultaneously. This suggests that the limit is dictated by how much a single model is loaded, not how much we are reading simultaneously. So if we split c1su2 into two models, we might be able to read more optics simultaneously, saving time and giving us the ability to measure for longer.

Attached are the results for all the core optics. Inferences will be made later in the week.

Note: Some measurements have very low coherence in IN2 channels in most of the damping frequency region, these loops need to be excited harder. (eg PIT, POS, YAW, on ITMs and ETMs).

 

Attachment 1: ALL_CORE_SUS_OLTF_2022-08-20-merged.pdf
ALL_CORE_SUS_OLTF_2022-08-20-merged.pdf ALL_CORE_SUS_OLTF_2022-08-20-merged.pdf ALL_CORE_SUS_OLTF_2022-08-20-merged.pdf ALL_CORE_SUS_OLTF_2022-08-20-merged.pdf ALL_CORE_SUS_OLTF_2022-08-20-merged.pdf ALL_CORE_SUS_OLTF_2022-08-20-merged.pdf ALL_CORE_SUS_OLTF_2022-08-20-merged.pdf ALL_CORE_SUS_OLTF_2022-08-20-merged.pdf
  17095   Fri Aug 19 15:36:10 2022 KojiUpdateGeneralSR785 C21593 CHA+ BNC broken

When Juan and I were working on the suspension measurement, I found that CHA didn't settle down well.

I inspected and found that CHA's + input seemed broken and physically flaky. For Juan's measurements, I plugged + channels (for CHA/B) and used - channels as an input. This seemed work but I wasn't sure the SR functioned as expected in terms of the noise level.

We need to inspect the inputs a bit more carefully and send it back to SRS if necessary.

How many SR785's do we have in the lab right now? And the measurement instruments like SR785 are still the heart of our lab, please be kind...

Attachment 1: PXL_20220819_195619620.jpg
PXL_20220819_195619620.jpg
Attachment 2: PXL_20220819_195643478.jpg
PXL_20220819_195643478.jpg
  17093   Fri Aug 19 15:20:14 2022 KojiUpdateGeneralNotice: SURF SUS test setup blocking the lab way

The setup was (at least partially) cleared.

Attachment 1: PXL_20220819_201318044.jpg
PXL_20220819_201318044.jpg
  17092   Fri Aug 19 14:46:32 2022 AnchalUpdateSUSOpen loop transfer function measurements for local damping loops of BHD optics

[Anchal, Tega]

As a first step to characterize all the local damping loops, we ran an open loop transfer function measurement test for all BHD optics, taking transfer function using band-limited (0.3 Hz to 10 Hz) gaussian noise injection at error points in different degrees of freedom. Plots are in the git repo. I'll make them lighter and post here.

We have also saved coherence of excitation at the IN1 test points of different degrees of freedom that may be later used to determine the cross-coupling in the system.

The test ran automatically using measSUSOLTF.py script. The script can run the test parallelly on all suspensions in principle, but not in practice because the cdsutils.getdata apparently has a limitation on how many real-time channels (we think it is 8 maximum) one can read simultaneously. We can get around this by defining these test points at DQ channels but that will probably upset the rtcds model as well. Maybe the thing to do is to separate the c1su2 model into two models handling 3 and 4 suspensions. But we are not sure if the limitation is due to fb or DAQ network (which will persist even if we reduce the number of testpoints on one model) or due to load on a single core of FE machines.

The data is measured and stored here. We can do periodic tests and update data here.


Next steps:

  • Run the test for old optics as well.
  • Fit the OLTF model with the measured data, and divide by the digital filter transfer function to obtain the plant transfer function for each loop.
  • Set maximum noise allowed in the local damping loop for each degree of freedom, and criteria for Q of the loop.
  • Adjust gains and or loop shape to reach the requirements on all the suspensions in a quantitative manner.
  • (optional) Add a BLRMS calculation stream in SUS models for monitoring loop performance and in-loop noise levels in the suspensions.
  • More frequency resolution, please. (KA)
  17091   Thu Aug 18 18:10:49 2022 KojiSummaryLSCFPMI Sensitivity

The overlapping plot of the calibrated error and control signals gives you an approximately good estimation of the freerun fluctuation, particularly when the open-loop gain G is much larger or much smaller than the unity.
However, when the G is close to the unity, they are both affected by "servo bump" and both signals do not represent the freerun fluctuation around that frequency.

To avoid this, the open-loop gain needs to be measured every time when the noise budget is calculated. In the beginning, it is necessary to measure the open-loop gain over a large frequency range so that you can refine your model. Once you gain sufficient confidence about the shape of the open-loop gain, you can just use measurement at a frequency and just adjust the gain variation (most of the cases it comes from the optical gain).

I am saying this because I once had a significant issue of (project-wide) incorrect sensitivity estimation by omitting this process.

  17090   Thu Aug 18 16:35:29 2022 CiciUpdateGeneralUGF linked to optical gain!

TL;DR: When the laser has good lock, the OLTF moves up and the UGF moves over!

-----------------------------------------------------------

Figured out with Paco yesterday that when the laser is locked but kind of weakly (mirrors on the optical table sliiightly out of alignment, for example), we would get a UGF around 5 kHz, but when we had a very strong lock (adjusting the mirrors until the spot was brightest) we would get a UGF around 13-17 kHz. Attached are some plots of us going back and forth (you can kind of tell from the coherence/error that the one with the lower UGF is more weakly locked, too). Error on the plots is propagated using the coherence data (see Bendat and Piersol, Random Data, Table 9.6 for the formula). 

-------------------------------------------------------------

Want to take data next week to quantitatively compare optical gain to UGF!

Attachment 1: rpi_OLG_2022_08_17_18_03_52.pdf
rpi_OLG_2022_08_17_18_03_52.pdf
Attachment 2: rpi_OLG_2022_08_17_18_00_50.pdf
rpi_OLG_2022_08_17_18_00_50.pdf
  17089   Thu Aug 18 14:49:35 2022 YehonathanSummaryLSCFPMI Sensitivity

{Yuta, Yehonathan}

We wrote a notebook found on Git/40m/measurements/LSC/FPMI/NoiseBudget/FPMISensitivity.ipynb for calculating the MICH, DARM (currently XARM), CARM (currently YARM) sensitivities in the FPMI lock which can be run daily.

The IN and OUT channels of each DOFs are measured at a certain GPS time and calibrated using the optical gains and actuation calibration measured in the previous post.

Attachment shows the results.

It seems like the UGFs for MICH and DARM (currently XARM) match the ones that were estimated previously (100Hz for MICH, 120Hz for DARM) except for CARM for which the UGF was estimated to be 250Hz and here seems to be > 1kHz.

Indeed one can also see that the picks in the CARM plot don't match that well. Calculation shows that at 250Hz OUT channel is 6 times more than the IN channel. Calibrations for CARM should be checked.

MICH sensitivity using REFL55 at high frequencies is not much better than what was measured with AS55.

DARM sensitivity at 10Hz is a factor of a few better than the single arm lock sensitivity.

Now it is time to do the budgeting.

Attachment 1: Sensitivity_Plot_1344133503.pdf
Sensitivity_Plot_1344133503.pdf
  17088   Wed Aug 17 11:10:51 2022 ranaUpdateComputersc1teststand rack mounting for CDS upgrade

we want to be able to run SimPlant on the teststand, test our new controls algorithms, test watchdogs, and any other software upgrades. Ideally in the steady state it will run some plants with suspensions and cavities and we will develop our measurement scripts on there also (e.g. IFOtest).

Quote:

[Tega, Yuta]

I keep getting confused about the purpose of the teststand. The view I am adopting going forward is its use as a platform for testing the compatibility of new hardware upgrade, instead of thinking of it as an independent system that works with old hardware.

  17087   Wed Aug 17 10:27:49 2022 CiciUpdateGeneralLocking X-arm AUX laser

TL;DR: Got the x-arm aux laser locked again and took more data - my fit on my transfer functions need improvement and my new method for finding coherence doesn't work so I went back to the first way! See attached file for an example of data runs with poor fits. First one has the questionable coherence data, second one has more logical coherence. (ignore the dashed lines.)

------------------------------------------------------------------------------------

  • The aux laser on the x-arm was still off after the power shutdown, so Paco and I turned it back on, and realigned the oplev of the ETMX - initial position was P = -0.0420, Y = -5.5391.
  • Locked the x-arm and took another few runs - was calculating coherence by I/Q demodulation of the buffers and then recombining the I/Q factors and then taking scipy.signal.coherence(), but for some reason this was giving me coherence values exclusively above 0.99, which seemed suspicious. When I calculated it the way I had before, by just taking s.s.coherence() of the buffers, I got a coherence around 1 except for in noisy areas of the data where it dropped more significantly, and seemed to be more correlated to the data. So I'll go back to using that way.
  • I also think my fits are not great - my standard error of the fits (calculated using the coherence as weight, see Table 9.6 of Random Data by Piersol and Bendat for the formula I'm using) are enormous. Now that I have a good idea that the UGF is between 1 - 15 kHz, I'm going to restrict my frequency band and try to fit just around where the UGF would be. 

--------------------------------------------------------------------------------

To do:

  • Reduce frequency band and take more data
  • Get fit with better standard error, use that error to calculate the uncertainty in the UGF!
Attachment 1: rpi_OLG_2022_08_16_17_00_41.pdf
rpi_OLG_2022_08_16_17_00_41.pdf
Attachment 2: rpi_OLG_2022_08_16_17_01_21.pdf
rpi_OLG_2022_08_16_17_01_21.pdf
  17086   Wed Aug 17 10:23:05 2022 TegaUpdateGeneralc1vac issues, pressure gauge replacement

- Disk full

I updated the configuration file '/etc/logrotate.d/rsyslog' to set a file sise limit of 50M on 'syslog' and 'daemon.log' since these are the two log files that capture caget & caput terminal outputs. I also reduce the number of backup files to 2.

controls@c1vac:~$ cat /etc/logrotate.d/rsyslog
/var/log/syslog
{
    rotate 2
    daily
    size 50M
    missingok
    notifempty
    delaycompress
    compress
    postrotate
        invoke-rc.d rsyslog rotate > /dev/null
    endscript
}

/var/log/mail.info
/var/log/mail.warn
/var/log/mail.err
/var/log/mail.log
/var/log/daemon.log
{
    rotate 2
    missingok
    notifempty
    size 50M
    compress
    delaycompress
    postrotate
        invoke-rc.d rsyslog rotate > /dev/null
    endscript
}
/var/log/kern.log
/var/log/auth.log
/var/log/user.log
/var/log/lpr.log
/var/log/cron.log
/var/log/debug
/var/log/messages
{
    rotate 4
    weekly
    missingok
    notifempty
    compress
    delaycompress
    sharedscripts
    postrotate
        invoke-rc.d rsyslog rotate > /dev/null
    endscript
}

- Vacuum gauge

The XGS-600 can handle 6 FRGs and we currently have 5 of them connected. Yes, having a spare would be good. I'll see about placing an order for these then.

Quote:

- Disk Full: Just use the usual /etc/logrotate thing

- Vacuum gauge

I rather feel not replacing P1a. We used to have Ps and CCs as they didn't cover the entire pressure range. However, this new FRG (=Full Range Gauge) does cover from 1atm to 4nTorr.

Why don't we have a couple of FRG spares, instead?

Questions to Tega: How many FRGs can our XGS-600 controller handle?

 

 

  17085   Wed Aug 17 07:35:48 2022 yutaBureaucracyGeneralMy wish list for IFO commissioning

FPMI related
- Better suspension damping HIGH
 - Investigate ITMX input matrix diagonalization (40m/16931)
 - Output matrix diagonalization
 * FPMI lock is not stable, only lasts a few minutes for so. MICH fringe is too fast; 5-10 fringes/sec in the evening.
- Noise budget HIGH
 - Calibrate error signals (actually already done with sensing matrix measurement 40m/17069)
 - Make a sensitivity curve using error and feedback signals (actuator calibration 40m/16978)
 * See if optical gain and actuation efficiency makes sense. REFL55 error signal amplitude is sensitive to cable connections.
- FPMI locking
 - Use CARM/DARM filters, not XARM/YARM filters
 - Remove FM4 belly
 - Automate lock acquisition procedure
- Initial alignment scheme
 - Investigate which suspension drifts much
 - Scheme compatible with BHD alignment
 * These days, we have to align almost from scratch every morning. Empirically, TT2 seems to recover LO alignment and PR2/3 seems to recover Yarm alignment (40m/17056). Xarm seems to be stable.
- ALS
 - Install alignment PZTs for Yarm
 - Restore ALS CARM and DARM
 * Green seems to be useful also for initial alignment of IR to see if arms drifted or not (40m/17056).
- ASS
 - Suspension output matrix diagonalization to minimize pitch-yaw coupling (current output matrix is pitch-yaw coupled 40m/16915)
 - Balance ITM and ETM actuation first so that ASS loops will be understandable (40m/17014)
- Suspension calibrations
 - Calibrate oplevs
 - Calibrate SUSPOS/PIT/YAW/SIDE signals (40m/16898)
 * We need better understanding of suspension motions. Also good for A2L noise budgeting.
- CARM servo with Common Mode Board
 - Do it with single arm first

BHD related
- Better suspension damping HIGH
 - Invesitage LO2 input matrix diagonalization (40m/16931)
 - Output matrix diagonalization (almost all new suspensions 40m/17073)
 * BHD fringe speed is too fast (~100 fringes/sec?), LO phase locking saturates (40m/17037).
- LO phase locking
 - With better suspensions
 - Measure open loop transfer function
 - Try dither lock with dithering LO or AS with MICH offset (single modulation)
 - Modify c1hpc/c1lsc so that it can modulate BS and do double demodulation, and try double demodulation
- Noise Budget HIGH
 - Calibrate MICH error signal and AS-LO fringe
 - Calibrate LO1, LO2, AS1, AS4 actuation using ITM single bounce - LO fringe
 - Check BHD DCPD signal chain (DCPD making negative output when fringes are too fast; 40m/17067)
 - Make a sensitivity curve using error and feedback signals
- AS-LO mode-matching 
 - Model what could be causing funny LO shape
 - Model if having low mode-matching is bad or not
 * Measured mode-matching of 56% sounds too low to explain with errors in mode-matching telescope (40m/16859, 40m/17067).

IMC related
- WFS loops too fast (40m/17061)
- Noise Budget
- Investigate MC3 damping (40m/17073)
- MC2 length control path

  17084   Wed Aug 17 01:18:54 2022 KojiUpdateGeneralNotice: SURF SUS test setup blocking the lab way

Juan and I built an analog setup to measure some transfer functions of the MOS suspension. The setup is blocking the lab way around the PD test bench.
Excuse us for the inconvenience. It will be removed/cleared by the end of the week.

Attachment 1: PXL_20220817_060428109.jpg
PXL_20220817_060428109.jpg
  17083   Tue Aug 16 18:22:59 2022 TegaUpdateComputersc1teststand rack mounting for CDS upgrade

[Tega, Yuta]

I keep getting confused about the purpose of the teststand. The view I am adopting going forward is its use as a platform for testing the compatibility of new hardware upgrade, instead of thinking of it as an independent system that works with old hardware.

The initial idea of clearing 1X7 cannot be done for now, because I missed the deadline for providing a detailed enough plan before Monday power up of the lab, so we are just going to go ahead and use the new rack as was initially intended and get the latest hardware and software tested here.

We mounted the DAQ, subnet and dolphin IX switches, see attachement 1. The mounting ears that came with the dolphin switch did not fit and so could not be used for mounting. We looked around the lab and decided to used one of the NavePoint mounting brackets which we found next to the teststand, see attachment 2.

We plan to move the new rack to the current location of the teststand and use the power connection from there. It is also closer to 1X7 so that moving the front-ends and switches to 1X7 should be straight forward after we complete all CDS upgrade testing.

Attachment 1: IMG_20220816_180157132.jpg
IMG_20220816_180157132.jpg
Attachment 2: IMG_20220816_175125874.jpg
IMG_20220816_175125874.jpg
  17082   Mon Aug 15 20:09:18 2022 KojiUpdateGeneralc1vac issues, 1 pressure gauge died

- Disk Full: Just use the usual /etc/logrotate thing

- Vacuum gauge

I rather feel not replacing P1a. We used to have Ps and CCs as they didn't cover the entire pressure range. However, this new FRG (=Full Range Gauge) does cover from 1atm to 4nTorr.

Why don't we have a couple of FRG spares, instead?

Questions to Tega: How many FRGs can our XGS-600 controller handle?

 

  17081   Mon Aug 15 18:06:07 2022 AnchalUpdateGeneralc1vac issues, 1 pressure gauge died

[Anchal, Paco, Tega]


Disk full issue:

c1vac was showing /var disk to be full. We moved all gunzipped backup logs to /home/controls/logBackUp. This emptied 36% of space on /var. Ideally, we need not log so much. Some solution needs to be found for reducing these log sizes or monitoring them for smart handling.


Pressure sensor malfunctioning:

We were unable to opel the PSL shuttter, due to the interlock with C1:Vac-P1a_pressure. We found that C1:Vac-P1a_pressure is not being written by serial_MKS937a service on c1vac. The issue was the the sensor itself has become bad and needs to be replaced. We believe that "L 0E-04" in the status (C1:Vac-P1a_status) message indicates a malfunctioning sensor.

Quick fix:

We removed writing of C1:Vac-P1a_pressure and C1:Vac-P1a_status from MKS937a and mvoed them to XGS600 which is using the sensor 1 from main volume. See this commit.

Now we are able to open PSL shutter. The sensor should be replaced ASAP and this commit can be reverted then.

  17080   Mon Aug 15 15:43:49 2022 AnchalUpdateGeneralComplete power shutdown and startup documentation

All steps taken have been recorded here:
https://wiki-40m.ligo.caltech.edu/Complete_power_shutdown_2022_08

  17079   Mon Aug 15 10:27:56 2022 KojiUpdateGeneralRecap of the additional measures for the outage prep

[Yuta Koji]

(Report on Aug 12, 2022)

We went around the lab for the final check. Here are the additional notes.

  • 1X9: The x-end frontend machine still had the AC power. The power strip to which the machine is connected was disconnected from the AC at the side of the rack. (Attachment 1)
  • 1X8: The vacuum rack still supplied the AC to c1vac. This was turned off at the UPS. (Attachment 2)
  • 1X6: VMI RFM hub still had the power. This was turned off at the rear switch. (Attachment 3)
  • PSL: The PSL door was open (reported above). Closed. (Attachment 4)
  • 1Y2: The LSC rack still had the DC power. The supplies were turned off at the KEPCO rack (the short rack). (Attachment 5)
    Note that the top-right supply for the +15V is not used. (The one in the empty slot got busted). We may need some attention to the left-most one in the second row. It indicated a negative current. Is this just the current meter problem or is the supply broken?
  • Control room: The CAD WS was turned off.

I declare that now we are ready for the power outage.

Attachment 1: PXL_20220812_234438097.jpg
PXL_20220812_234438097.jpg
Attachment 2: PXL_20220812_234655309.jpg
PXL_20220812_234655309.jpg
Attachment 3: PXL_20220812_234748559.jpg
PXL_20220812_234748559.jpg
Attachment 4: rn_image_picker_lib_temp_b5f3e38d-796c-4816-bc0e-b11ba3316cbe.jpg
rn_image_picker_lib_temp_b5f3e38d-796c-4816-bc0e-b11ba3316cbe.jpg
Attachment 5: PXL_20220812_235429314.jpg
PXL_20220812_235429314.jpg
  17078   Fri Aug 12 13:40:36 2022 JCUpdateGeneralPreparing for Shutdown on Saturday, Aug 13

[Yehonathan, JC]

Our first step in preparing for the Shutdown was to center all the OpLevs. Next is to prepare the Vacuum System for the shutdown.

 

  17077   Fri Aug 12 02:02:31 2022 KojiUpdateGeneralPower Outage Prep: nodus /home/export backup

Took the backup (snapshot) of /home/export as of Aug 12, 2022

controls@nodus> cd /cvs/cds/caltech/nodus_backup
controls@nodus> rsync -ah --progress --delete /home/export ./export_220812 >rsync.log&

As the last backup was just a month ago (July 8), rsync finished quickly (~2min).

  17076   Thu Aug 11 17:15:33 2022 CiciUpdateGeneralMeasuring AUX Laser UGF with Red Pitaya

TL;DR: Have successfully measured the UGF of the AUX laser on my Red Pitaya! Attached is one of my data runs (pdf + txt file). 

---------------------------------------------------------------

  • Figured out how to get a rudimentary coherence (use scipy.signal.coherence to get Cxy = abs(Pxy)**2/(Pxx*Pyy), then find what point is the closest to the frequency I'm inserting on that iteration of the swept sine and get the coherence closest to that). Not precisely the coherence at the frequency I'm inserting though, so not perfect... more of a lower bound of coherence.
  • Figured out how to get the UGF from the data automatically (no error propagation yet... necessary next step)
  • Put my red pitaya in the X-arm AUX laser control electronics (thank you to Anchal for help figuring out where to put it and locking the x-arm.) Counts dropped from 4500 to 1900 with the x-arm locked, so 58% mode matching. I lose lock at an amplitude >0.05 or so.
  • Wrote a little script to take data and return a time-stamped text file with all the parameters saved and a time-stamped pdf of the TF magnitude, UGF, phase, and coherence, so should be easy to take more data next time!

----------------------------------------------------------------

  • need to take more accurate coherence data
  • need to propagate uncertainty on UGF (probably high...)
  • take more data with higher coherence (the file attached doesn't have great coherence and even that was one of my better runs, will probably increase averaging since increasing amplitude was a problem)
Attachment 1: rpi_OLG_2022_08_11_16_51_53.pdf
rpi_OLG_2022_08_11_16_51_53.pdf
Attachment 2: rpi_OLG_2022_08_11_16_51_53.txt
# frequency start: 500.0
# frequency stop: 50000.0
# samples: 50
# amplitude: 0.01
# cycles: 500
# max fs: 125000000.0
# N: 16384UGF: 9264.899326705621
# Frequency[Hz] Magnitude[V/V] Phase[rad] Coherence
4.999999999999999432e+02 5.216612299292965105e+01 -7.738468629291910261e-01 7.660920305860696722e-02
5.492705709937790743e+02 3.622076363933444298e+01 -5.897393740774580229e-01 3.183076012979469405e-01
... 49 more lines ...
  17075   Thu Aug 11 16:48:59 2022 ranaUpdateComputer Scripts / ProgramsNDS2 updates

We had several problems with our NDS2 server configuration. It runs on megatron, but I think it may have had issues since perhaps not everyone was aware of it running there.

  1. channel lists were supposed to updated regularly, but the nds2_nightly script did not exist in the specified directory. I have moved it from Joe Areeda's personal directory (/home/nds2mgr/joework/server/src/utils/) to nds2mgr/channel-tracker/.
  2. The channel history files (/home/nds2mgr/channel-tracker/channel_history/) are stored on the local megatron disk. These files had grown up to ~50 GB over tha past several years. I backed these up to /users/rana/, and then wiped them out so that the NDS could regen them. Now that the megatron local disk is not full, it seems to work in giving raw data.
  3. Need to confirm that this serves up trend data (second and minute)
  4. I think there is a nds2-server package for Debian, so we should update megatrons OS to the preferred flavour of DebIan and use that. Who to get to help in this install?

Since Megatron is currently running the "Shanghai" Quad-core Opteron processor from ~2009,  its ~time to replace it with a more up to date thing. I'll check with Neo to see if he has any old LDAS leftovers that are better.

  17074   Wed Aug 10 20:51:14 2022 TegaUpdateComputersCDS upgrade Front-end machine setup

Here is a summary of what needs doing following the chat with Jamie today.

 

Jamie brought over the KVM switch shown in the attachment and I tested all 16 ports and 7 cables and can confirm that they all work as expected.

 

TODO

1. Do a rack space budget to get a clear picture of how many front-ends we can fit into the new rack

2. Look into what needs doing and how much effort would be needed to clear rack 1X7 and use that instead of the new rack. The power down on Friday would present a good opportunity to do this work on Monday, so get the info ready before then. 

3. Start mounting front-ends, KVM and dolphin network switch

4. Add the BOX rack layout to the CDS upgrade page.

Attachment 1: IMG_20220810_171002928.jpg
IMG_20220810_171002928.jpg
Attachment 2: IMG_20220810_171019633.jpg
IMG_20220810_171019633.jpg
  17073   Wed Aug 10 20:30:54 2022 TegaSummarySUSCharacterisation of suspension damping

[Yuta, Tega]

We diagnosed the suspension damping of the IMC/BHD/recycling optics by kicking the various degree of freedom (dof) and then tuning the gain so that we get a residual Q of approx. 5 in the cases where this can be achieved.

MC1: Good
MC2: SIDE-YAW coupling, but OK
MC3: Too much coupling between dofs, NEEDS ATTENTION
LO1: Good
LO2: Good
AS1: POS-PIT coupling, close to oscillation, cnt2um off, NEEDS ATTENTION
AS4: PIT-YAW coupling, cannot increase YAW gain because of coupling, No cnt2um, No Cheby, NEEDS ATTENTION
PR2: No cnt2um, No Cheby
PR3: POS-PIT coupling, cannot increase POS/PIT/YAW gain because of coupling, No cnt2um, No Cheby, NEEDS ATTENTION
SR2: No cnt2um

Attachment 1: BHD_SUSPIT_KICK.png
BHD_SUSPIT_KICK.png
Attachment 2: BHD_SUSPOS_KICK.png
BHD_SUSPOS_KICK.png
Attachment 3: BHD_SUSSIDE_KICK.png
BHD_SUSSIDE_KICK.png
Attachment 4: BHD_SUSYAW_KICK.png
BHD_SUSYAW_KICK.png
Attachment 5: IMC_SUSPIT_KICK.png
IMC_SUSPIT_KICK.png
Attachment 6: IMC_SUSPOS_KICK.png
IMC_SUSPOS_KICK.png
Attachment 7: IMC_SUSSIDE_KICK.png
IMC_SUSSIDE_KICK.png
Attachment 8: IMC_SUSYAW_KICK.png
IMC_SUSYAW_KICK.png
Attachment 9: PRSR_SUSPIT_KICK.png
PRSR_SUSPIT_KICK.png
Attachment 10: PRSR_SUSPOS_KICK.png
PRSR_SUSPOS_KICK.png
Attachment 11: PRSR_SUSSIDE_KICK.png
PRSR_SUSSIDE_KICK.png
Attachment 12: PRSR_SUSYAW_KICK.png
PRSR_SUSYAW_KICK.png
  17072   Wed Aug 10 19:36:45 2022 KojiBureaucracyGeneralLab cleaning and discovery

During the cleaning today, we found many legacy lab items. Here are some policies what should be kept / what should be disposed

Dispose

  • VME crates and VME electronics as long as they are not in use
  • Eurocard SUS modules that are not in use.
  • Eurocard crates (until we remove the last Eurocard module from the lab)
  • Giant steel plate/palette (like a fork lift palette) along the Y arm. (Attachment 1)
  • An overhead projector unit.

Keep

  • Spare Eurocard crates / ISC/PZT Eurocard modules
  • Boxes of old 40m logbooks behind the Y arm (see Attachment 2/3).
  • Ink-plotter time-series data (paper rolls) of 1996 IFO locking (Attachment 4). Now stored in a logbook box.
  • A/V type remnants: Video tapes / video cameras / casette tapes as long as they hold some information in it. i.e. Blank tapes/blank paper rolls can be disposed.
Attachment 1: steel_plate.jpg
steel_plate.jpg
Attachment 2: logbook1.jpg
logbook1.jpg
Attachment 3: logbook2.jpg
logbook2.jpg
Attachment 4: paper_plots.jpg
paper_plots.jpg
  17071   Wed Aug 10 19:24:19 2022 ranaUpdateGeneralWorking Red Pitaya VNA

                 Boom!

 

  17070   Wed Aug 10 15:33:59 2022 CiciUpdateGeneralWorking Red Pitaya VNA

TL;DR: I am now able to inject a swept sine and measure a transfer function with python on my Red Pitaya! Attached is a Bode plot for a swept sine from 1 - 30 MHz, going through a band pass filter of 9.5 - 11.5 MHz.

------------------------------------------------------------------------------------

  • Spent too long trying to get pyRPL to work, do not recommend. The code on their website has a lot of problems (like syntax-error level problems), and is ultimately designed to open up and start a GUI, which is not what I want even if it did work.
  • Found some code on the git repository of someone at Delft University of Technology, worked better but still not great (oscilloscope/spectrum analyzer functions were alright, but couldn't successfully run a VNA with it, and overcomplicated). Helped me figure out appropriate decimation factors. Realized it was not using the FPGA to get TF data but instead just collecting a lot of time trace data and then taking an FFT in the code to get the TF, which wasn't ideal.
  • Eventually switched to using the Red Pitaya SCPI server to talk to the Red Pitaya myself, successful! I inject a swept sine with a for loop that just cycles through frequencies and takes the transfer function at each one.
    • Was originally getting the transfer function by using scipy.signal.csd() and scipy.signal.welch() to get Pxy and Pxx and dividing, and then just finding the closest point in the frequency spectrum to the frequency I was inserting.
    • Switched to doing IQ demodulation myself: where x(t) is the measurement before the band pass filter and y(t) is the measurement after, taking the mean of (x(t) * cos(2pi*freq)) = a1, mean(x*sin()) = a2, mean(y*cos()) = b1, mean(y*sin()) = b2, and then TF(freq) = (b1 + i b2)/(a1 + i a2).
    • Unfortunately still taking time trace data and then calculating the TF instead of using the FPGA, but I have not found anything online indicating that people are able to get VNA capabilities on the Red Pitaya without collecting and sending all the time trace data... I'm still not sure if that's actually a Red Pitaya capability yet.

-------------------------------------------------------------------------------------

To do:

  • Will go take measurements of the AUX laser loop with the RPi! Have a good diagram of when I did it with the SR785 so it shouldn't be too hard hopefully.
  • Figure out how to get coherence data!!
  • Figure out how to get the RPi on the wifi. Right now I'm just plugging the RPi into my computer. Paco and I were working on this before and had trouble finding old passwords... Hopefully will not be too much of a roadblock.
Attachment 1: rpi_vna_test.pdf
rpi_vna_test.pdf
  17069   Tue Aug 9 19:54:31 2022 yutaSummaryLSCFPMI locking tonight

[Tega, Anchal, Yuta]

We resored FPMI locking settings. Below is the summary of locking configurations tonight.
To ease the lock acquisition, the step to feedback POX11_I to ETMX and POY11_I to MC2 before POX and POY mixing was necessary tonight.

CARM (YARM):
 - 0.5 * POX11_I + 0.5 * POY11_I handed to 0.5 * REFL55_I
 - YARM filter module, FM4,5 for acquisition, FM1,2,3,6,8 triggered, C1:LSC-YARM_GAIN = 0.012
 - Actuation on -0.77 * MC2
 - UGF ~ 250 Hz

DARM (XARM):
 - 0.5 * POX11_I - 0.5 * POY11_I handed to 4.6 * AS55_Q (it was 2.5 in 40m/17012)
 - XARM filter module, FM5 for acquisition (no FM4), FM1,2,3,6,8 triggered, C1:LSC-XARM_GAIN = 0.015
 - Actuation on 0.5 * ETMX - 0.5 * ETMY
 - UGF ~ 120 Hz

MICH:
 - 1 * REFL55_Q (turned on after XARM and YARM acquisition)
 - MICH filter module, FM4,5,8 for acquisition, FM2,3 triggered, C1:LSC-MICH_GAIN = +40
 - Actuation on 0.5 * BS
 - UGF ~ 100 Hz

Measured sensing matrix:
Sensing Matrix with the following demodulation phases
{'AS55': 200.41785156862835, 'REFL55': 93.7514468401475, 'POX11': 105.08325063571438, 'POY11': -11.343909976281823}
Sensors              DARM                    CARM                   MICH
C1:LSC-AS55_I_ERR_DQ 5.27e-02 (-154.105 deg) 2.83e-01 (132.395 deg) 1.17e-04 (-40.1051 deg)
C1:LSC-AS55_Q_ERR_DQ 3.99e-02 (-151.048 deg) 1.42e-02 (125.504 deg) 1.41e-04 (-2.42846 deg)
C1:LSC-REFL55_I_ERR_DQ 5.59e-02 (77.6871 deg) 1.15e+00 (-44.589 deg) 3.55e-04 (69.2585 deg)
C1:LSC-REFL55_Q_ERR_DQ 1.84e-03 (16.3186 deg) 3.35e-03 (125.67 deg) 4.59e-05 (4.18718 deg)
C1:LSC-POX11_I_ERR_DQ 1.54e-01 (-157.852 deg) 6.07e-01 (-42.1078 deg) 5.55e-05 (73.3963 deg)
C1:LSC-POX11_Q_ERR_DQ 6.83e-05 (-148.591 deg) 6.37e-04 (121.983 deg) 1.35e-06 (43.7201 deg)
C1:LSC-POY11_I_ERR_DQ 1.85e-01 (36.1624 deg) 5.73e-01 (-43.1776 deg) 2.12e-04 (82.16 deg)
C1:LSC-POY11_Q_ERR_DQ 2.16e-05 (130.937 deg) 6.38e-05 (-173.194 deg) 1.40e-06 (47.5416 deg)

FPMI locked periods:
  - 1344129143 - 1344129520
  - 1344131106 - 1344131305
  - 1344133503 - 1344134020

Next:
- Restore CM servo for CARM

ELOG V3.1.3-